Bank uk 25. - U VU b) $200 (6) $400 d) $600 Customer deposits & $1000, deconto at Bank of Canada - If the Bank of Canada tranplease answer all questions 25-29! the information needed is written above but if you can't read it properly it's listed as follows: in an imperial bank, customer deposits = $1000, Deposits at the bank of Canada = $10, cash on hand = $60, and desired reserve coefficient = 5%

Thanks so much!
Answer #1

25) When government of Canada deposit money into bank of Canada,there was two entries in the balance sheet of bank ba Canada.

First Increase in liability by same amount of deposit.

Second Increase in assest by same amount ,enter as reserves.

When bank of Canada deposit some of that deposit into reliable bank ,the assest doesn't Increase instead,the reserve Decreases and loan Increases.and total assest remind unchanged.

Option c is right

26) The deposit ,for reliable bank is liabilities ( thet have to pay interest),so liabilities Increases.

But those deposit can be used as lending loans from which they earn interst,so assest also Increases.

So option A is right.

27) multiplier=1/ mps=1/0.2=5

∆y=Multiplier*∆G

∆G=∆Y/5

∆Y=200

∆G=200/5=40

Option A is correct

28)Real wage is actual wage when CPI =100

So actual wage in 2018.

120/100=actual wage 2018/10

Actual wage=12.

Wage increase by 5% in 2019

Wage in 2019=12*1.05=12.6

Option C is right

29) using rule of 70,

70/6=1.667{ time take to double)

(1.06)^n=30,000/15000

(1.06)^11.895661046=2

n=11.895661046